uestioN in out Given Tio)o. P(n)- gる 0 ou
Answer #1

P( 0in ) = 0.1 ; P( 1in ) = 0.5 ;  P( 2in ) = 0.4

P(0out | 0in ) = 0.1   P(0out | 1in ) = 0.3 P(0out | 2in ) = 0.1

P(1out | 0in ) = 0.8   P(1out | 1in ) = 0.3 P(1out | 2in ) =

P(2out | 0in ) =   P(2out | 1in ) = P(2out | 2in ) = 0.75

A. 1. P(2out | 0in ) = 1 - ( P(0out | 0in ) + P(1out | 0in ) ) = 1 - (0.1 + 0.8) = 1 - 0.9 = 0.1

2. P(2out | 1in ) = 1- ( P(0out | 1in ) + P(1out | 1in ) ) = 1 - ( 0.3 + 0.3) = 1 - 0.6 = 0.4

3. P(1out | 2in ) = 1 - (P(0out | 2in ) +   P(2out | 2in )) = 1 - (0.1 +0.75) = 1 - 0.85 = 0.15

B. P(1out) = P(1out | 0in ) * P(0in) +  P(1out | 1in ) * P(1in) + P(1out | 2in ) * P(2in)

= 0.8 * 0.1 + 0.3*0.5 + 0.15*0.4

= 0.08 + 0.15 + 0.06 = 0.29

C. P(0out) = P(0out | 0in ) * P(0in) +  P(0out | 1in ) * P(1in) + P(0out | 2in ) * P(2in)

= 0.1 *0.1 + 0.3 *0.5 + 0.1*0.4

= 0.01 + 0.15 + 0.04 = 0.2

P(2in U 0out ) = P(2in | 0out ) * P( 0out ) = P(0out | 2in ) * P(2in)

Therefore,

P(2in | 0out ) =  \tiny \frac{P(0out | 2in ) * P(2in) }{P(0out) }

= \tiny \frac{0.1 * 0.4 }{0.2 } = 0.04 / 0.2 = 0.2

D. P(1in | 0out ) =  \tiny \frac{P(0out | 1in ) * P(1in) }{P(0out) }

= \tiny \frac{0.3 * 0.5 }{0.2 } = 0.15/ 0.2 = 0.75